For hydrogen, an electron is in 2p state. Obtain the magnitude of orbital angular momentum and possible



z-components of orbital angular momentum 𝐿⃗ . What would be the possible values of total angular



momentums (𝐽 ) and associated magnetic moments 𝜇𝐽



⃗⃗⃗

1
Expert's answer
2022-01-10T09:06:48-0500

The magnitude of orbital angular momentum is


"L=\dfrac h{2\pi}\sqrt{l(l+1)}."


The z-components are


"L_z=m_l\dfrac h{2\pi}."